Solve for the initial pressure in lb/ft² (before heated) and the heat added in Btu if a 20 m³ steel tank confines 200 lbs of certain gas consisting of 6% Hgas of Cp=3.419 btu/lb°R, R=766.54 ft.lb/lb°R, and 94% Sgasof Cp=0.15413 kJ/kg°K, R=2.7196 ft.lb/lb°R is being heated at constant volume until the absolute temperature doubles when initially at 12°C.


1
Expert's answer
2021-11-05T01:13:03-0400

Initial absolute temperature

"= 12+273.15=285.15K"


Final absolute temperature

"=2*285.15=570.3K"


Converting units given in other systems to FPS system:


"285.15K=53.6°F\quad 570.3K=566.87°F"